Grade 12th passAlgebra

Solve the following equations.
3/7 + x = 17/7
Solve the following equations.
3/7 + x = 17/7

Harshit Singh
5 Years agoGrade 12th pass

2 Answers

Harshit Singh
5 Years ago
Dear Student

3/7+x=17/7
x = 17/7–3/7
x=14/7
x=2

Thanks
Shraddha Singh
5 Years ago
Dear Student,
 
We can solve it as follows:
3/7 + X = 17/7
X = 17/7 - 3/7
X = (17-3)/7
X = 14/7
X = 2
 
Hope it helps!
Thank you ✌️